(X^2) y"-2(x^2+x) y'+(x^2+2x+2) y=0
1
Expert's answer
2021-09-16T00:46:27-0400

Solution;

Assume;

"y=\displaystyle\sum_{n=0}^{\infin}a_nx^{r+n}"

"y'=\displaystyle\sum_{n=0}^{\infin}a_n(r+n)x^{r+n-1}"

"y''=\displaystyle\sum_{n=0}^{\infin}a_n(r+n)(r+n-1)x^{r+n-2}"

Substitute y,y' and y'' into the given equation to obtain;

"\displaystyle\sum_{n=0}^{\infin}a_n(r+n)(r+n-1)x^{r+n}-\displaystyle\sum_{n=0}^{\infin}2a_n(r+n)x^{r+n+1}-\displaystyle\sum_{n=0}^{\infin}2a_n(r+n)x^{r+n}+\displaystyle\sum_{n=0}^{\infin}a_nx^{r+n+2}+\displaystyle\sum_{n=0}^{\infin}2a_nx^{r+n+1} +\displaystyle\sum_{n=0}^{\infin}2a_nx^{r+n}"

We shift the terms as follows;

"\displaystyle\sum_{n=0}^{\infin}a_n(r+n)(r+n-1)x^{r+n} -\displaystyle\sum_{n=1}^{\infin}2a_{n-1}(r+n)x^{r+n} -\displaystyle\sum_{n=0}^{\infin}2a_n(r+n)x^{r+n}+\displaystyle\sum_{n=2}^{\infin}a_{n-2}x^{r+n}+\displaystyle\sum_{n=1}^{\infin}2a_{n-1}x^{r+n}+\displaystyle\sum_{n=0}^{\infin}2a_nx^{r+n}"

At n=0;

"a_0[r(r-1)-2r+2]=0"

But "a_0\neq0"

Hence;

"r^2-3r+2=0"

"(r-2)(r-1)=0"

Therefore;

r=2 or 1

At n=1;

"a_1(r+1)r-2a_0(r+1)-2a_1(r+1)+2a_0+2a_1=0"

"a_1r^2+a_1r-2a_0r-2a_0-2a_1r-2a_1+2a_0+2a_1=0"

Simplify as;

"a_1(r^2+r-2r)-2a_0r=0"

"a_1=\frac{2a_0r}{r(r-1)}"

To solve ,we take r=2;

"a_1=2a_0"

At n=2 ,all terms produces,hence we have a general formula;

"a_n(r+n)(r+n-1)-2a_{n-1}(r+n)-2a_n(r+n)+a_{n-2}+2a_{n-1}+2a_n=0"

Combine as follows;

"a_n[(2+n)(n+1)-2(2+n)+2]=2a_{n-1}[2+n+2]+a_{n-2}"

Hence;

"a_n=\frac{2a_{n-1}(4+n)+a_{n-2}}{(2+n)(n+1)-2(n+2)+2}"

This is the recurrence relationship using r=2;

Hence we find other terms;

n=2;

"a_2=\frac{2a_1(6)+a_0}{4.3-8+2}" ="\frac{12a_1+a_0}{6}"

Since "a_1=2a_0" ,by substitution;

"a_2=\frac{25a_0}{6}"

When n=3;

"a_3=\frac{2a_2(8)+a_1}{5.4-10+2}=\frac{14a_2+a_1}{12}"

By substitution ;

"a_3=\frac{181a_0}{36}"

When n=4;

"a_4=\frac{2a_3(8)+a_2}{6.5-16+2}=\frac{16a_3+a_2}{20}"

By substitution;

"a_4=\frac{1523}{360}a_0"

The solution of the equation becomes;

"y=x^2\displaystyle\sum_{n=0}^{\infin}a_nx^{n}"

Hence ;

"y=x^2[a_0+2a_0x+\frac{25a_0}{6}x^2+\frac{181}{36}a_0x^3+\frac{1523}{360}a_0x^4+..."